OpenStudy (anonymous):

Two trains leave stations 504 miles apart at the same time and travel toward each other. One train travels at 100 miles per hour while the other travels at 80 miles per hour. How long will it take for the two trains to meet? Do not do any rounding.

OpenStudy (anonymous):

You did not write any numbers

OpenStudy (anonymous):

Do you see numbers now ?

OpenStudy (anonymous):

hint: Both will meet at the same time at some point. so... speed=distance/time or time=dist/speed. since time is equal for both, only speed and distance are different so: time=distance/speed of one train =distance/speed of the 2nd train

OpenStudy (anonymous):

I want you to understand the idea and fill inthe numbers, give it a try

OpenStudy (anonymous):

the distance for one would be x and the other would be (504-x)

OpenStudy (anonymous):

x/100=(504-x)/80

OpenStudy (anonymous):

once you know x you can find the time the train going at 100 would take to reach that distance.

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

I didn't get it >.<

OpenStudy (anonymous):

The two trains are 504 miles apart and start at the same time at different speeds

OpenStudy (anonymous):

Now they will meet at some point at the same time, so the even though one will run faster and travel more distance than the other but since we want to know when they meet so the time it take both of them to get to that point is the same

OpenStudy (anonymous):

do you get this far

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so now this time is diatance/ speed, right?. since speed=distance/time got it?

OpenStudy (anonymous):

yes i got it but how can i now that the distance for one would be x and the other would be (504-x) ?

OpenStudy (anonymous):

Because the total distance between them is 504 miles and we don't know what is the distance at which they meet. So if say train A going at 100m/h travels x miles the other travels 504-x miles.

OpenStudy (anonymous):

The remaining distance

OpenStudy (anonymous):

Since the time is the same for both dist/speed of trainA=dist/speed of trian B So x/100= (504-x)/80 and we get x=280 , they meet at a point which is 280 miles from where the train going at 100m/hr started out. In other words t= 280/100 or 2.8 hours . The other train going travels 504-280=224 miles. Just to check if we are right 224/80=2.8 hours again. I can explain again if you didn't get.

OpenStudy (anonymous):

Okay i got it ! .. thank you so much =)